RC meeting and Run Plan, Monday November 8th
RC Meeting Information
RC Daily Meetings are at 3:15pm, remote connection link: https://bluejeans.com/256205934
Accelerator Status
The accelerator is delivering beam to Halls C and D.
Hall A Status and News
We’re still waiting for signed OSPs to move the spectrometers. The magnet test group is coming out tomorrow to look at repairs. Tomorrow afternoon will have the final walkthrough. From the moment we get the OSP we likely have 2 days of equipment moving and 1 day of testing. Therefore it is likely that we will start taking beam at the end of the week. The exit beam line was bled up and the target chamber vacuum seems to have held. Dave Meekins will be away this week so Chris Keith will condense the target when the time comes. SBS 7 is essential so if there are extended delays some other setting will have to be truncated. Shifts are cancelled until Wednesday morning.
Subsystem Updates
- HCAL: N/A
- BBCal: We had a presentation showing that the gain balancing work was very successful. When the magnets are in their new positions more cosmic data will be needed to see the effect of the field on the gain. About 1 hour of data is sufficient.
- Timing hodoscope: No major updates. First full round of calibrations are now done.
- Grinch: Scalers turned down to 10 Hz. More voltage scan will be done this week.
- GEMS: Testing different solutions to the high current draw issue. Some new voltage divider boards at UVa which could go in early December. The new UV plane is ready in the Test Lab when Jessie is able to install it. Jack has mixing system that is being tested at the moment.
- Beamline: N/A
- DAQ: Running most of the time. Some issues with MPDs getting stuck.
- Replay analysis: Getting replay running on counting house machines. Comparing GEM occupancies to simulation both raw and with cuts.